OpenStudy (unklerhaukus):

\[\sum_{n=1}^{∞}1/n!\]

OpenStudy (anonymous):

We know that:\[e^x=\sum_{n=0}^\infty \frac{x^n}{n!}\] Letting x=1 gives: \[e^1=\sum_{n=0}^{\infty}\frac{1^n}{n!}\] Therefore: \[e-1=\sum_{n=1}^{\infty}\frac{1}{n!}\]

OpenStudy (unklerhaukus):

so about 1 .718

OpenStudy (anonymous):

yep, thats right.

OpenStudy (unklerhaukus):

thanks

OpenStudy (unklerhaukus):

\[so, -∑(iπ)^n/n!=1 ?\]

OpenStudy (anonymous):

yes that is correct lolol